OpenStudy (anonymous):

The point (1,2) has the distance ________ from the line defined by y=−x3+1. could somebody solve this problem

Mehek (mehek14):

is the line \(\Large{y = -x^3+1}\)

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

do you know the differential calculus?

OpenStudy (anonymous):

no its my precalculus lecture's problem

OpenStudy (michele_laino):

ok! Then I try to search for another method, please wait...

OpenStudy (amistre64):

unless its a line, i don think there is a geometric way to approach it

OpenStudy (amistre64):

d^2 = (x-k)^2 + (y-k)^2 is minimal unless that setup simplifies to something simple ...

OpenStudy (amistre64):

x-h not x-k

OpenStudy (anonymous):

http://upload.wikimedia.org/math/c/a/e/cae82210ba48deb8ff6fb0134a90b36e.png i think this is the formula to solve it but when i solve it was wrong

OpenStudy (amistre64):

just wondering d^2 = (x-1)^2 + (-x^3+1-2)^2 d^2 = x^2 -2x +1 + x^6+1+2x^3 d^2 = x^6 +2x^3 +x^2 -2x +2 unless the right side is a perfect square, i dont think this has a chance ...

OpenStudy (amistre64):

where does your formula come from?

OpenStudy (anonymous):

instructor

OpenStudy (amistre64):

if all you are doing is trying to memorize formulas, then you are not being taught anything ... computer programs can determine the distances based on formulas

OpenStudy (anonymous):

its not what i am doing i just suggest it

OpenStudy (amistre64):

d^2 = (x-h)^2 + (y-k)^2 taking the derivative 2d d' = 2x' (x-h) + 2y' (y-k) d d' = (x-h) + y' (y-k) (x-h) + y' (y-k) d' = ----------- = 0 when (x-h) + y'(y-k) = 0 d

OpenStudy (anonymous):

i do not know calculus yet it an precalculus question

OpenStudy (amistre64):

this is how i would approach it since its not a line

OpenStudy (michele_laino):

I found another method, nevertheless we have to know the meaning of the slope of a tangent line to a function as the first derivative of that function.

OpenStudy (amistre64):

good luck

OpenStudy (michele_laino):

do you know the concept of first derivative of a function?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok! Now let's consider a generic point Q on your function y=-x^3+1, namely \[Q = \left( {{x_0},{y_0}} \right)\]

OpenStudy (michele_laino):

or: \[Q = \left( {{x_0},{y_0}} \right) = \left( {{x_0}, - x_0^3 + 1} \right)\]

OpenStudy (michele_laino):

is it ok?

OpenStudy (michele_laino):

Q belongs to our function f(x) = -x^3+1

OpenStudy (anonymous):

yes i am ok with this solution

OpenStudy (michele_laino):

ok! Now I have to write the tangent line, to our function, at that point. The slope m, of that tangent line is: \[m = f'\left( x \right) = - 3x_0^2\]

OpenStudy (michele_laino):

so the equation of our tangent line, is: \[y - \left( { - x_0^3 + 1} \right) = - 3x_0^2\left( {x - {x_0}} \right)\]

OpenStudy (michele_laino):

please tell me, when I may continue

OpenStudy (michele_laino):